A pipeline connecting two tanks contains four standard elbows, a plug valve that is fully open and a gate valve that is half open. The line is commercial steel pipe, 25 mm internal diameter, length 120 m. The properties of the fluid are: viscosity 0.99 mNM_2 s, density 998 kg/m3. Calculate the total pressure drop due to friction when the flow rate is 3500 kg/h


1
Expert's answer
2021-02-01T09:30:04-0500

Velocity of flow V = Q/A = (3500 kg / (3600 s x 998 kg/m3)) / ("\pi"/4 x (0.025 m)2) = 1.986 m/s

Reynolds number, Re = "\rho"VD/"\mu" = (998 kg/m3 x 1.986 m/s x 0.025 m) / 0.099Ns/m2 = 500

The Reynolds number is less than 2000. So the flow is laminar.

For laminar flow, Friction factor f = 64/Re = 64/500 = 0.128

Head loss, HL = f(L/D)(V2/2g) = 0.128 x (120 m / 0.025 m) x ((1.986 m/s)2 / 2 x 9.81 m/s2) = 121.4 m

Pressure drop = "\gamma" HL = 998 kg/m3 x 9.81 m/s2 x 121.4 m = 11.9 x 105 Ns/m2 = 12 bar